Get Help

Online: Request Help
Phone: (707) 826-4357 • Hours
Email: help@humboldt [dot] edu
Walk-In: Library 101 • Hours
Reset HSU Password
System Status

Services For

System Status Archive

System Issue start Who is affected What happened
Other Oct 5 2015 - 08:00 Some people trying to access campus systems
Wireless Network Oct 8 2015 - 11:00 Some people trying to access the HSU network via HSUWireless-Secure or HSUWireless ITS is investigating. *Update 1 - TNS has identified a ARP resource limitation in the router that serves both HSUWireless and HSUWireless-Secure. Configuration changes will be applied over night to resolve the issue.* *Update 2 - Configuration change has been applied, ARP limitation has been resolved.TNS will monitor.*
CMS/Peoplesoft Oct 7 2015 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window, the following items will be installed: MP 17.04 and MP 18.00.
Other Oct 6 2015 - 05:30 All Users The Database Server Portal depends (Poplar) was found unstable, and had to be cold-booted. Portal services will need to be restarted. ITS is investigating. Update: Portal is once again available.
Campus Data Network Oct 5 2015 - 09:45 Many campus network users, wired and wireless TNS is investigating. Check back for updates.
Other Oct 1 2015 - 04:00 Anyone using the Guest Wireless and Resnet networks Planned upgrade for CNI core refresh project.
Computer Labs Sep 30 2015 - 18:10 All users of AutoCAD 2012/2015, Maya 2015, Mathematica 9/10, MATLAB (Mathematics labs only), Riverware, NI LabVIEW, SolidWorks, Codon Code Aligner, and Trimble GPS Pathfinder Office The Mathematica license server service and license needed updating.
Computer Labs Sep 30 2015 - 17:20 All users of AutoCAD 2012/2015, Maya 2015, Mathematica 9/10, MATLAB (Mathematics labs only), Riverware, NI LabVIEW, SolidWorks, Codon Code Aligner, and Trimble GPS Pathfinder Office The Mathematica license server service and license needed updating.
Pharos Pay-for-Print Sep 27 2015 - 16:23 ** Services are once again available. ** Anyone trying to print to campus public computers. ** Services are once again available. ** ITS is currently investigating. Will update as more details emerge.
Other Sep 26 2015 - 22:00 Trinity users One of the datastores used by the virtual machine ran out of space while backing up. This caused a complete crash of the server. We are working to resolve the issue. Update: Maintenance is taking longer than expected. Update2: We are unable to get a firm ETA on recovery. It will hopefully be done in the morning. Update3: Server should be available
Off-Campus Internet Connection Sep 27 2015 - 05:00 Campus network users/services accessing the internet Border Firewalls are experiencing an issue that prevents redundancy from working as intended. Vendor recommends a reload. ****UPDATE*** Reboot completed, services are restored.
Other Sep 21 2015 - 12:00 ** ITS is still investigating ** Administrative Affairs Users ** ITS is still investigating ** ITS is looking into the problem, check back for updates; UPDATE #1: we are working to move the data to a new server, hopefully completed by 8pm Sunday, 9/27/2015 8pm. Then we will follow data integrity comparison and copy over any missing data. Should be resolved by Monday, 9/28/2015 at 8am.
Pharos Pay-for-Print Sep 23 2015 - 14:20 People printing to Pharos Pay-for-Print printers from Linux computers We are currently investigating the cause which has not yet been determined. Update to follow. UPDATE:issue has been resolved
Other Sep 22 2015 - 07:00 Anyone accessing voicemail Devices need to be migrated to a new network switch in preparation for the final phase of the CNI network upgrade project.
Other Sep 21 2015 - 09:10 Some users logging on or trying to access files ITS is looking into the problem UPDATE 1: The systems have returned to normal functionality.
Campus Data Network Sep 19 2015 - 00:01 Campus network users/services accessing the internet Need to reload firewalls to accommodate additional network modules.
Other Sep 17 2015 - 17:30 Some users. VMware has dropped support for ESXi 4 in their api, and as a result we are unable to backup virtual machines on our ESXi cluster. We are migrating to ESXi 5 as soon as possible. Previous migrations have gone smoothly, and we do not expect issues. We expect that the final migrations will be finished during the monthly maintenance window on Sept. 20th.
Other Sep 20 2015 - 05:00 All users Scheduled Monthly Maintenance.
CMS/Peoplesoft Sep 16 2015 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window the following items will be installed: HUM_SDRC_FIX_004 modification.
Other Sep 15 2015 - 20:00 Nolij users We need to add disk space to the nolij datastore before it runs out. Update: The maintenance is taking much longer than expected. The resolution ETA has been revised accordingly. Update: Maintenance has finished.
Other Sep 15 2015 - 17:30 Anyone using the CSU Portal, Zimride, and other Shibboleth service providers Added new service provider to shibboleth idp required restart of service.
Other Sep 13 2015 - 04:00 See services below The virtual systems need to be shutdown to be migrated to a new virtual infrastructure.
Other Sep 8 2015 - 17:30 All users VMware has dropped support for ESXi 4 in their api, and as a result we are unable to backup virtual machines on our ESXi cluster. We are migrating to ESXi 5 as soon as possible. We will be migrating more servers this week as scheduling permits, and will be issuing more systats as the migration progresses. UPDATE- 9/8/2015 7:15PM: Migrations scheduled for this evening are complete.
Other Sep 8 2015 - 17:30 Shibboleth Services including OBI and CSYOU Portal. Added new Shibboleth Service Provider.
Campus Phone System Sep 3 2015 - 10:18 On campus phone users trying to make off campus calls It appears that AT&T is having an outage somewhere along our phone route. TNS is working with AT&T. Will update the system when we know more. Update #1 - there is still no known time for resolution Update #2 AT&T has restored service about 6am this morning---
Computer Labs Sep 2 2015 - 12:15 Users of ENVI, ENVI+IDL and Vlab versions of JAWS and Magic The ENVI+IDL trial license was updated and the server was rebooted to confirm the service would come back up after a reboot.
Computer Labs Sep 2 2015 - 12:15 Users of ENVI, ENVI+IDL and Vlab versions of JAWS and Magic The ENVI+IDL trial license is expired this week and needs to be updated.
CMS/Peoplesoft Sep 2 2015 - 18:00 Anyone who uses Student Center, Faculty Center, Human Resources, Student Records, Student Financials, Financial Aid, Admissions During our regularly scheduled maintenance window the following items will be installed: MP 17.03, SDRC fix 003 and Admissions Student Center App Status fix 2.
Other Aug 26 2015 - 08:27 Users of Data Warehouse, IRP, DARS, Nolij, APEX apps, and Account Settings (possibly others) There is a hardware fault involving the sanstore which is causing a degradation in IO capability which becomes noticeable during heavy IO, specifically during nightly processing. The vendor is scheduled to come out on Friday and repair the issue. It is advised that until the vendor comes out any large jobs that can be put on hold should be (especially those that run during the night). This will allow for jobs that must run every night to have a higher chance to succeed in a timely manner as well as reduce that chance for database overload resulting in a loss of service. Update: Vendor has completed service and initial testing of the database systems appear to have returned to normal state. Nightly processes will be running again tonight and data should be up to date for tomorrow.
Other Aug 24 2015 - 16:40 Students and instructors using TurningPoint Clickers Students and instructors using TurningPoint Clickers are seeing errors at the system level with Clickers not functioning correctly. We will likely need to upgrade the plugin on the server and will provide updates as we learn more. UPDATE 2: We are continuing to work with Turning Technologies this morning on this issue. UPDATE 3: The issues with Turning Technologies Clickers has been fixed. Users will need to re-register their clickers in Moodle. We apologize for the inconvenience.
feedback